# Cooperative Task Spaces for Multi-Arm Manipulation Control based on Similarity Transformations

### v1

@@ IV-B2 Circle Reaching
The last scenario for the cooperative circle features a manipulator and a humanoid that are collaborating for the reaching of the desired circle, as shown in Figure 10(a). Although, we use a humanoid in this example, it could be easily replaced by a human in order to model a human-robot collaboration scenario. Since this extension provides an interesting opportunity for future work, we discuss it further in Section V-E. The bivector command trajectory can be seen in Figure 10(b).
@@ IV-C Geometric Nullspace
Using $\bm{\dot{q}}_{d,p}$, the first robot still reaches the target point, while staying on the circle. We show this in Figure 15. It can be seen, that by executing the motion in the nullspace of the cooperative task space, which corresponds to the cooperative geometric primitive, i.e. the circle, the first robot follows the circle. Therefore, the similarity distance is zero during the task.
@@
IV-D Teleoperation Demonstration
@@ V Discussion
Our unified control framework based on cooperative similarity transformations, generalizes the rigid body transformation of a single robot’s end-effector to the cooperative control of multiple kinematic chains. The experimental results demonstrated this approach across a variety of robotic systems, ranging from a bi-manual setup to scenarios involving collaborating humanoids. Given that, the framework is purely geometric, and considering the variety of systems explored, the kinematic simulations show both the strengths and limitations of our method, which are further discussed in the following subsections.
@@ V-A Connection to Traditional Single-Arm Control
Throughout this article, we have mentioned several times that the cooperative similarity transformation and its Jacobians fulfill the same role as the rigid body transformation to the robot end-effector that is used in single-arm systems. Since similarity transformations form a seven-dimensional manifold, as opposed to the six-dimensional on of rigid body transformations, the modeling using the cooperative similarity transformations only introduces one additional dimension to achieve cooperative control behaviors of highly complex robotic systems. Although the involved mathematics that we derived in this article might seem unfamiliar, all the findings in the literature for traditional control methods of single-arm systems remain valid and are directly applicable to the scenarios shown in this article. We have shown this via differential kinematics control in the teleoperation experiments in Section IV-D. This further highlights how the cooperative modeling greatly facilitates the control of complex robotic systems.
@@ V-B Geometric Nullspace
The systems considered in our experiments range from a bimanual setup with 14 degrees of freedom (DoF) (see Figure 18(a)) to collaborating humanoids with 34 DoF in total. The resulting systemss are highly redundant, and thus are inherently challenging to control. Even when restricted to the task space, the dimensionality remains high, i.e. ranging from 12 to 24 DoF. Notably, encoding the cooperative task space using geometric primitives significantly reduces the effective dimensionality of the problem. The similarity transformations associated with these primitives have at most 7 DoF, accounting for translation, rotation, and uniform scaling. Depending on the symmetry of the primitive, the effective number of DoF can be even lower. For instance, a sphere (see Figure 12(a)) admits only 4 DoF, as rotation does not alter its configuration. As can be inferred from the teleoperation experiment given in Figure 16, this reduction in dimensionality is critical: no teleoperation device exists for 28-DoF systems, nor can a human or an existing algorithm feasibly control such a high-dimensional system directly. In contrast, using our proposed method for controlling the translation and dilation of a target sphere is straightforward.
@@
A second key aspect of the redundancy resolution arises from the use of geometric nullspaces. Geometric primitives naturally define subspaces of the underlying algebra, giving rise to geometric nullspace formulations for control. This structure ensures that the controller is inherently stiff when the system deviates from the primitive, and compliant when moving along it, as we previously showed in the context of optimal control for a single robotic arm [22]. This principle extends directly to the cooperative geometric primitives introduced in this work, as illustrated in Figures 18(a) and 18(b).
The geometric nullspaces induced by these primitives enable a decoupling of orthogonal control objectives in a coordinate-free manner, while eliminating the need for basis-specific representations or manual tuning of off-diagonal matrix terms. For example, aligning the end-effector with a target line does not constrain motion along the line, enabling the controlled application of contact forces in that direction. We previously demonstrated this feature with a single robot in an impedance control application on curved surfaces [4]. As shown in Figure 18(a), this property naturally extends to cooperative settings, such as coordinated object lifting.
@@ V-G Different Types of Geometric Primitives
In this work, we described the general cooperative similarity task spaces for multiple parallel kinematic chains. We wanted to emphasize the usage of the cooperative similarity transformation as an equivalent to the end-effector pose. Thus, in our control and optimization examples, we always assumed the target geometric primitives to be of the same type as the task space geometry, as listed in Table I. In general, however, it is possible to formulate task objectives and constraints involving different types of geometric primitives. In principle, this could be approached via the outer product minimization that we presented for a single manipulator in [22], where cooperative geometric primitive could then be used to reach a desired point. Geometric relationships, such as intersections between primitives, can also be fully described in the algebra, and thus can be used in optimization problems. Note that these relationships are well-defined, i.e. they will always yield a mathematically valid result even if the primitives do not intersect.
V-H Practical Considerations

### v2

@@ IV-B2 Circle Reaching
The last scenario for the cooperative circle features a manipulator and a humanoid that are collaborating for the reaching of the desired circle, as shown in Figure 10(a). Although we use a humanoid in this example, it could be easily replaced by a human in order to model a human-robot collaboration scenario. Since this extension provides an interesting opportunity for future work, we discuss it further in Section V-E. The bivector command trajectory can be seen in Figure 10(b).
@@ IV-B4 Sphere Reaching
IV-B5 Line Constraint
The cooperative geometric primitives can not only serve as task objectives but also as constraints for secondary tasks. In the previous examples, we demonstrated reaching problems in which the cooperative primitive transitions from one configuration to another. Here, we illustrate how the cooperative line primitive can be used to impose geometric constraints on the motion of a dual-arm system, while simultaneously executing individual end-effector tasks.
We use the Unitree G1 humanoid robot with its two arms. Together with the waist joints, it provides 17 degrees of freedom. The cooperative line primitive is formed by the two wrist points. In the first scenario, shown in Figure 14(a), we task the right hand to reach a target point, while the left hand remains stationary. This represents an asymmetric reaching task in which only one arm is actively controlled, while the other provides a fixed reference. The cooperative line connecting the two wrists therefore changes freely as the right hand moves towards its target, which can be observed from the trajectory of the line parameters in Figure 14(b).
In the second scenario, we introduce an additional constraint on the cooperative line, which must now remains horizontal throughout the motion. As shown in Figure 15(a), the right hand again reaches a target point, but now the left hand must move in coordination to maintain the horizontal orientation of the line primitive. This coordinated motion ensures that the line parameter corresponding to the vertical component remains constant, which is reflected in Figure 15(b).
This example demonstrates the versatility of the cooperative geometric primitive framework for encoding both task objectives and geometric constraints. By specifying constraints on the cooperative primitive, such as maintaining a horizontal orientation, the framework automatically generates coordinated motion across multiple kinematic chains. This capability is particularly useful for tasks like carrying objects with specific orientation requirements, or maintaining formation constraints during collaborative manipulation.
(a) The initial configuration of the robot is shown in white and the final one in gray. The right hand moves (shown with red trajectory) to reach a target point while the left hand remains stationary. The blue line represents the cooperative line primitive connecting the two wrists.
(b) Trajectory of the line parameters over time. The parameters change as the right hand moves while the left hand remains fixed, showing unconstrained evolution of the cooperative line primitive.
Figure 14: Asymmetric reaching task with a cooperative line primitive. The right hand reaches a target point while the left hand stays in place, allowing the cooperative line to change freely.
(a) The initial configuration of the robot is shown in white and the final one in gray. Both hands move (shown with red trajectories) in a coordinated manner to maintain the horizontal orientation of the cooperative line primitive while the right hand reaches the target.
(b) Trajectory of the line parameters over time. Compared to the unconstrained case, the value of $\bm{e}_{03i}$ remains constant, reflecting the horizontal orientation constraint on the cooperative line primitive.
Figure 15: Constrained reaching task with a cooperative line primitive. The right hand reaches a target point while the cooperative line is constrained to remain horizontal, requiring coordinated motion of both arms.
@@ IV-C Geometric Nullspace
Using $\bm{\dot{q}}_{d,p}$, the first robot still reaches the target point, while staying on the circle. We show this in Figure 17. It can be seen that, by executing the motion in the nullspace of the cooperative task space, which corresponds to the cooperative geometric primitive, i.e. the circle, the first robot follows the circle. Therefore, the similarity distance is zero during the task.
@@
IV-D Geometric Singularity
In addition to the classical kinematic singularities that affect individual manipulators, cooperative systems exhibit geometric singularities that arise from the degeneration of the cooperative geometric primitive itself. These singularities occur when the geometric primitive loses its defining characteristics, such as when three points forming a circle become collinear, or when a circle’s radius approaches zero or infinity.
To illustrate this phenomenon, we consider the same setup as in Section IV-B2 (i.e., three Franka robots with a total of 21 degrees of freedom), where the cooperative geometric primitive is a circle defined by the three end-effector points. We task the system with reaching a target configuration that is a geometric singularity, as shown in Figure 18. In this configuration, the three end-effector points approach a collinear arrangement, causing the circle to degenerate. Note that we are not controlling the cooperative circle here, but instead track its evolution while one end-effector reaches the line formed by the other two.
The trajectory of the circle parameters over time is shown in Figure 19(a). As the system approaches the singular configuration, several of the circle parameters drop rapidly towards zero, indicating the degeneration of the circle primitive. Correspondingly, the bivector trajectory in Figure 19(b) shows a drastic change in the cooperative similarity transformation. These large components reflect that lines in CGA can essentially be seen as circles with infinite radius (i.e. the similarity transformation becomes infinite).
The effect of the geometric singularity on the system’s manipulability is depicted in Figure 19(c). As expected, several components of the task space manipulability decrease significantly when approaching the singularity, with some dropping to zero values. This loss of manipulability indicates that certain directions in the cooperative task space become difficult or impossible to control, which is the property of a singular configuration.
This example demonstrates that geometric singularities present an important consideration for the control of cooperative systems modeled by geometric primitives. Unlike classical kinematic singularities that depend solely on individual joint configurations, geometric singularities are intrinsic to the cooperative task space representation and can occur even when all individual manipulators are far from their kinematic singularities. Strategies for singularity avoidance in this context would need to monitor the geometric properties of the cooperative primitive and potentially modify the desired trajectory or task space representation when approaching degenerate configurations.
Figure 18: Configuration of the three Franka robots approaching a geometric singularity. The end-effector points approach a collinear arrangement, causing the cooperative circle primitive (shown in blue) to degenerate to a line (shown in red). The initial robot configurations are shown with lighter color.
(a) Trajectory of the circle primitive parameters over time. The parameters drop rapidly to zero as the circle degenerates when the three points approach collinearity.
(b) Trajectory of the similarity transformation bivector. Several components diverge, reflecting the large transformation for a line, i.e. a circle with infinite radius.
(c) Eigenvalues of the task space manipulability over time. Several components drop to zero values, indicating loss of controllability in certain task space directions due to the geometric singularity.
Figure 19: Effect of geometric singularity in a cooperative system of three Franka robots. The system passes through a configuration where the circle primitive degenerates as the three end-effector points become nearly collinear.
IV-E Teleoperation Demonstration
@@ V Discussion
Our unified control framework based on cooperative similarity transformations generalizes the rigid body transformation of a single robot’s end-effector to the cooperative control of multiple kinematic chains. The experimental results demonstrated this approach across a variety of robotic systems, ranging from a bi-manual setup to scenarios involving collaborating humanoids. Given that the framework is purely geometric, and considering the variety of systems explored, the kinematic simulations show both the strengths and limitations of our method, which are further discussed in the following subsections.
@@ V-A Connection to Traditional Single-Arm Control
Throughout this article, we have mentioned several times that the cooperative similarity transformation and its Jacobians fulfill the same role as the rigid body transformation to the robot end-effector that is used in single-arm systems. Since similarity transformations form a seven-dimensional manifold, as opposed to the six-dimensional one of rigid body transformations, the modeling using the cooperative similarity transformations only introduces one additional dimension to achieve cooperative control behaviors of highly complex robotic systems. Although the involved mathematics that we derived in this article might seem unfamiliar, all the findings in the literature for traditional control methods of single-arm systems remain valid and are directly applicable to the scenarios shown in this article. We have shown this via differential kinematics control in the teleoperation experiments in Section IV-E. This further highlights how the cooperative modeling greatly facilitates the control of complex robotic systems.
@@ V-B Geometric Nullspace
The systems considered in our experiments range from a bimanual setup with 14 degrees of freedom (DoF) (see Figure 22(a)) to collaborating humanoids with 34 DoF in total. The resulting systems are highly redundant, and thus are inherently challenging to control. Even when restricted to the task space, the dimensionality remains high, i.e. ranging from 12 to 24 DoF. Notably, encoding the cooperative task space using geometric primitives significantly reduces the effective dimensionality of the problem. The similarity transformations associated with these primitives have at most 7 DoF, accounting for translation, rotation, and uniform scaling. Depending on the symmetry of the primitive, the effective number of DoF can be even lower. For instance, a sphere (see Figure 12(a)) admits only 4 DoF, as rotation does not alter its configuration. As can be inferred from the teleoperation experiment given in Figure 20, this reduction in dimensionality is critical: no teleoperation device exists for 28-DoF systems, nor can a human or an existing algorithm feasibly control such a high-dimensional system directly. In contrast, using our proposed method for controlling the translation and dilation of a target sphere is straightforward.
@@
A second key aspect of the redundancy resolution arises from the use of geometric nullspaces. Geometric primitives naturally define subspaces of the underlying algebra, giving rise to geometric nullspace formulations for control. This structure ensures that the controller is inherently stiff when the system deviates from the primitive, and compliant when moving along it, as we previously showed in the context of optimal control for a single robotic arm [undefu]. This principle extends directly to the cooperative geometric primitives introduced in this work, as illustrated in Figures 22(a) and 22(b).
The geometric nullspaces induced by these primitives enable a decoupling of orthogonal control objectives in a coordinate-free manner, while eliminating the need for basis-specific representations or manual tuning of off-diagonal matrix terms. For example, aligning the end-effector with a target line does not constrain motion along the line, enabling the controlled application of contact forces in that direction. We previously demonstrated this feature with a single robot in an impedance control application on curved surfaces [undefc]. As shown in Figure 22(a), this property naturally extends to cooperative settings, such as coordinated object lifting.
@@ V-C Singularities of Cooperative Primitives
V-C1 Conditions for Well-Definedness
The cooperative geometric primitive $X_{c}(\bm{q})$ defined in Equation (14) is well-defined if and only if the outer product of the defining points yields a non-degenerate geometric entity. This requirement imposes geometric constraints on the relative positions of the end-effector points. Specifically:
  • Pointpair/Line: Two points must be distinct, i.e., $P_{1}\neq P_{2}$. When the points coincide, the pointpair degenerates to a single point.
  • Circle/Plane: Three points must be non-collinear. When the three points lie on a straight line, the circle degenerates to a line (or equivalently, a circle with infinite radius in CGA).
  • Sphere: Four points must be non-coplanar. When all four points lie in the same plane, the sphere degenerates to a plane.
As the system approaches these degenerate configurations, the cooperative geometric primitive becomes increasingly ill-defined. In the limiting case, the primitive cannot uniquely determine the cooperative similarity transformation $V_{Sc}(\bm{q})$, and the control objective loses its geometric meaning.
V-C2 Rank Deficiency and Jacobian Conditioning
The cooperative Jacobian matrices derived in Section III-D maintain sufficient rank as long as the cooperative primitive remains well-defined and the individual manipulator Jacobians are not singular. However, near degenerate configurations, the cooperative Jacobian exhibits poor conditioning and may lose rank.
Consider the example of a cooperative circle approaching collinearity as shown in Section IV-D. As the three defining points align, the radius of the circle grows unbounded, and certain components of the similarity transformation become ambiguous. Consequently, the geometric similarity Jacobian $\bm{\mathcal{J}}^{G}_{Sc}(\bm{q})$ grows unbounded in specific directions, similar to the behavior observed near kinematic singularities in single-arm systems.
[…]
V-C3 Practical Implications and Mitigation Strategies
In practice, the limitations discussed above can be mitigated through singularity avoidance strategies using the manipulability measure from Section III-K to detect and avoid degenerate configurations, similar to standard techniques in single-arm control. While these challenges are inherent to the cooperative primitive framework, they are not fundamentally different from the analogous challenges faced in traditional multi-robot coordination or single-arm control near singularities. The key distinction is that the cooperative geometric primitive provides an explicit, coordinate-free characterization of the degenerate configurations, making them easier to detect and avoid.
@@ V-G Different Types of Geometric Primitives
In this work, we described the general cooperative similarity task spaces for multiple parallel kinematic chains. We wanted to emphasize the usage of the cooperative similarity transformation as an equivalent to the end-effector pose. Thus, in our control and optimization examples, we always assumed the target geometric primitives to be of the same type as the task space geometry, as listed in Table I. In general, however, it is possible to formulate task objectives and constraints involving different types of geometric primitives. In principle, this could be approached via the outer product minimization that we presented for a single manipulator in [undefu], where cooperative geometric primitive could then be used to reach a desired point. Geometric relationships, such as intersections between primitives, can also be fully described in the algebra, and thus can be used in optimization problems. Note that these relationships are well-defined, i.e. they will always yield a mathematically valid result even if the primitives do not intersect.
V-H Numerical Analysis
A critical numerical consideration in the implementation of our framework concerns the conditioning of the logarithmic map used to extract similarity bivectors from versors. The logarithmic maps for rotors, translators, dilators, and similarity transformations presented in Appendix LABEL:sec:transformation_groups_in_cga all involve trigonometric functions in their denominators. The primary source of numerical instability arises when the rotation angle is a whole multiple of $\pi$, which manifests as the scalar part of the rotor $\langle V_{R}\rangle_{0}$ approaching $\pm 1$. In these configurations, the denominator $\sin(\cos^{-1}(\langle V_{R}\rangle_{0}))$ in Equation (LABEL:eq:log_rotor) becomes zero, leading to division by zero. These singular cases are straightforward to detect by monitoring the scalar component $|\langle V_{R}\rangle_{0}|$ of the rotor before computing the logarithm. Note that rotors are essentially identical to quaternions, and therefore the same challenges and solutions to the conditioning of the logarithmic map apply. Similar considerations apply to the other transformation groups, though the rotation component typically presents the most significant numerical challenge.
V-I Computational Efficiency
A similarity bivector $B_{S}\in\mathbb{B}_{S}$ requires storage of 7 scalar components, while the corresponding similarity transformation versor $V_{S}\in S$ requires 12 components (1 scalar, 7 bivector, and 4 quadvector components). The geometric product of two similarity transformations requires 112 floating-point multiplications and 100 additions, according to the geometric product rules where certain blade combinations are known to result in zero. For a single manipulator, a homogeneous transformation matrix in SE(3) maps the end effector pose to the target pose through a single matrix product. Exploiting the zeros in the last row of the homogeneous matrices, this multiplication requires about 39 multiplications and 34 additions. Consequently, for three or more independent kinematic chains, the total number of operations required by similarity transforms or by separate rigid body transformations similar in terms of computational complexity. Considering the fact that the similarity transformations encode up to four kinematic chains, this computational overhead is acceptable, since it in turns enables more efficient constraint evaluation on all kinematic chains simultaneously. An in-depth comparison of dual quaternions (i.e. the rigid transformation part of the similarity transformations) versus transformation matrices can be found in [dantamRobustEfficientForward2021].
V-J Practical Considerations